### v1

@@ Abstract
The acquisition cost for large, annotated motion datasets remains a critical bottleneck for skeletal-based Human Activity Recognition (HAR). Although Text-to-Motion (T2M) generative models offer a compelling, scalable source of synthetic data, their training objectives, which emphasize general artistic motion, and dataset structures fundamentally differ from HAR’s requirements for kinematically precise, class-discriminative actions. This disparity creates a significant domain gap, making generalist T2M models ill-equipped for generating motions suitable for HAR classifiers. To address this challenge, we propose KineMIC (Kinetic Mining In Context), a transfer learning framework for few-shot action synthesis. KineMIC adapts a T2M diffusion model to an HAR domain by hypothesizing that semantic correspondences in the text encoding space can provide soft supervision for kinematic distillation. We operationalize this via a kinetic mining strategy that leverages CLIP text embeddings to establish correspondences between sparse HAR labels and T2M source data. This process guides fine-tuning, transforming the generalist T2M backbone into a specialized few-shot Action-to-Motion generator. We validate KineMIC using HumanML3D as the source T2M dataset and a subset of NTU RGB+D 120 as the target HAR domain, randomly selecting just 10 samples per action class. Our approach generates significantly more coherent motions, providing a robust data augmentation source that delivers a +23.1% accuracy points improvement. Animated illustrations and supplementary materials are available at https://lucazzola.github.io/publications/kinemic.
@@ 1 Introduction
Human Activity Recognition (HAR) has become a cornerstone in a multitude of fields, including sports performance analysis, human-robot collaboration, and intelligent surveillance [Dentamaro2024HumanAR]. Skeletal-based HAR remains a fundamental modality, frequently employed due to its lightweight representation, robustness to environmental variations, and inherently privacy-preserving nature [yan2018spatialtemporalgraphconvolutional]. However, the performance of deep learning models for this task is fundamentally limited by the availability of large-scale, accurately annotated datasets. The acquisition and precise labeling of such high-quality, task-specific data is notoriously expensive and labor-intensive, creating a significant bottleneck that hampers progress, particularly in few-shot settings [Dentamaro2024HumanAR, HungCuong2023DeepLF]. This core challenge of data scarcity in HAR is what we aim to address.
@@
To mitigate this fundamental data bottleneck, much of modern few-shot recognition research focuses on applying strategies such as meta-learning and metric-learning frameworks directly to the action recognition model [wanyan2025comprehensivereviewfewshotaction]. An alternative approach leverages generative models to create novel synthetic samples, thereby augmenting the small training set [fukushi2024fewshot]. When it comes to motion synthesis, current research is dominated by Text-to-Motion (T2M) synthesis models, due to the growing interest in text as a conditioning modality. This interest has driven the creation and collection of large-scale T2M datasets [t2m_humanml, mahmood2019amassarchivemotioncapture], which has, in turn, led to the development of powerful deep models [guo2023momaskgenerativemaskedmodeling, t2m_humanml, tevet2022humanmotiondiffusionmodel] that can be employed as general motion priors for other generative tasks [sawdayee2025dancelikechickenlowrank, shafir2023humanmotiondiffusiongenerative, tevet2024closdclosingloopsimulation]. T2M priors are strategically appealing as a generative solution because the scalability of their annotation, using free-form, descriptive text, is significantly easier to achieve for general diversity than collecting high-volume, kinematically specific data for HAR.
@@
While the community moves towards developing foundation models for 3D humans [khirodkar2024sapiensfoundationhumanvision, wang2025prompthmrpromptablehumanmesh], the challenge of effectively exploiting such general T2M priors for specialized, downstream tasks like HAR remains underexplored. Our work concentrates on this challenge: adapting a general T2M prior to function as an Action-to-Motion (A2M) synthetic data generator for a specific target HAR domain. This transformation is non-trivial due to a significant domain gap characterized by two key factors. First, a semantic discrepancy exists where source T2M data uses descriptive text, while target HAR requires generation based on discrete action labels. Second, a kinematic gap exists between the broad, fluid motions of the source domain and the short, atomic motions required for HAR. The pre-trained T2M model, being a generalist, is therefore ill-equipped to meet requirements for reliable HAR classification.
@@
To address these challenges and bridge this domain gap for few-shot action synthesis, we propose KineMIC (Kinetic Mining In Context, Fig. 1). Our method employs teacher-student architecture, wherein a frozen teacher, pre-trained on a source T2M dataset, guides the fine-tuning of a student model on the limited target HAR domain. Our method starts by establishing a semantic correspondence between the sparse target action labels and the rich source textual descriptions through CLIP [radford2021learningtransferablevisualmodels] text encoder. Secondly, relevant motion sub-sequences, extracted from the vast source dataset, are used to turn the general-purpose student into a specialized generator. The main contributions of this work are as follows:
@@
  • We introduce KineMIC, a teacher-student framework, which ultimately adapts a general T2M diffusion prior [tevet2022humanmotiondiffusionmodel] to a specific HAR domain with minimal data, proving its effectiveness at improving HAR accuracy.
@@ 2.1 Skeletal-based HAR
Recognizing human activity from skeletal data is a pivotal research area in computer vision. Early approaches modeled motion as a time series using RNNs and LSTMs [du2015hierarchical, shahroudy2016nturgbdlargescale]. A paradigm shift occurred with the introduction of Graph Convolutional Networks (GCNs) [yan2018spatialtemporalgraphconvolutional], which re-conceptualized the skeleton as a graph to model spatio-temporal dependencies. This led to rapid advancements using adaptive graph structures [shi2019twostreamadaptivegraphconvolutional], refined GCNs [chen2021channelwisetopologyrefinementgraph, chi2022infogcn, duan2022pysklgoodpracticesskeleton, liu2020disentanglingunifyinggraphconvolutions], and more recently, specialized architectures like 3D convolutional networks [duan2022revisitingskeletonbasedactionrecognition] and Transformers [do2024skateformerskeletaltemporaltransformerhuman, plizzari2021spatio]. Despite these advancements, the field still struggles with data scarcity, as the high capacity of modern models leads to severe overfitting in limited-data regimes. Our work addresses this bottleneck by proposing a novel synthesis framework leveraging deep T2M models, to create the necessary training data, supporting HAR applications in few-shot settings.
@@ 2.2 Generative Models for 3D Skeleton-based Motion
Realistic human motion synthesis evolved from early VAEs and GANs conditioned on discrete actions [Guo_2020, petrovich2021actionconditioned3dhumanmotion] to highly expressive models. This progress was fueled by large motion capture datasets [t2m_humanml, mahmood2019amassarchivemotioncapture] and richer conditioning signals, advancing from text [petrovich2022temosgeneratingdiversehuman] to even music [zhang2025motionanythingmotiongeneration] as a modality. The current state-of-the-art is dominated by Denoising Diffusion Probabilistic Models (DDPMs) [tevet2022humanmotiondiffusionmodel, zhao2023modiffactionconditioned3dmotion] and masked generative models [guo2023momaskgenerativemaskedmodeling], which achieve impressive generative fidelity. However, these T2M models, trained for general character animation, inherently lack the kinematic specificity required for downstream HAR tasks. Our proposed KineMIC framework addresses this domain gap by using a kinetic mining strategy to distill HAR-relevant knowledge from the T2M backbone, enabling its application as a specialized data generator.
@@ 2.3 Few-Shot HAR with Generative Models
The high cost of annotated motion data bottlenecks Few-Shot Human Activity Recognition (FSHAR). While classifier-side methods (e.g., metric/meta-learning) [wanyan2025comprehensivereviewfewshotaction] are the most popular approach, they are inherently constrained by the limited kinematic diversity of the few-shot support set. A less-explored alternative is deep generative data augmentation [leng2025scalinghumanactivityrecognition, lupion2024dataaugmentation]. Most notably, recently Fukushi et al. [fukushi2024fewshot] demonstrated this using a GAN with cross-domain regularization. To bypass GAN notorious training instability and explore a new avenue, we propose leveraging modern T2M diffusion models for few-shot motion synthesis. In direct contrast to [fukushi2024fewshot], our approach employs a "semantics-first" matching strategy. By utilizing CLIP [radford2021learningtransferablevisualmodels] text encoding space, we disentangle the synthesis process from the kinematic limitations of the few-shot set.
@@ 3 Problem Formulation
The core challenge we address is the adaptation of a pre-trained T2M generative model for data augmentation within the context of few-shot HAR. Our goal is to leverage the extensive kinematic knowledge contained in a large source (i.e. prior) domain to synthesize a high volume of diverse, class-specific motion sequences for a target domain, thereby enhancing the performance of a downstream HAR classifier. Let a skeletal motion sequence be defined as $\mathbf{x}=\{x(j)\in\mathbb{R}^{d}\}_{j=1}^{n}$, where $n$ is the number of frames and $d$ is the dimensionality of the pose representation. We consider two distinct domains:
@@ 4 Methodology
Figure 3: (Left) Core pipeline: dashed lines show detached gradients; dotted lines show weight sharing. MIC receives frame-wise tokens from teacher ($G^{P}$) and student ($G^{T}$) streams, producing latents $z^{P}$ and $z^{\bar{T}}$ for contrastive alignment and window mining. (Right) MIC module: attention-enhanced biGRU [bahdanau2016neuralmachinetranslationjointly, gru] encoder aggregating frame-wise motion tokens into context-aware latents.
@@ 4.1 Teacher-Student Architecture
KineMIC builds on two identical MDM [tevet2022humanmotiondiffusionmodel] models: a frozen teacher $G^{P}$ and a trainable student $G^{T}$. Both models retain MDM’s core structure and diffusion paradigm: predicting clean motion $\mathbf{x}_{0}$ from noised input $\mathbf{x}_{t}$ under a conditioning signal $c$. We refer to [tevet2022humanmotiondiffusionmodel] for full diffusion process details. To preserve prior knowledge, $G^{T}$ is fine-tuned with Low-Rank Adaptation (LoRA) [hu2021loralowrankadaptationlarge], which updates only a small fraction of weights through low-rank matrices while keeping the vast majority of pre-trained parameters frozen. Furthermore, while the student inherits the frozen text embedding layer used in T2M, we introduce a learnable action embedding (Fig. 2) that enables $G^{T}$ specialization to discrete HAR labels.
@@ 4.2 Semantic Alignment via Soft Positive Search
Target labels in $\mathcal{D}^{T}$ lack the descriptive richness of T2M captions in $\mathcal{D}^{P}$, creating a one-to-many mapping (single label $\rightarrow$ multiple text prompts). We bridge this via CLIP-based [radford2021learningtransferablevisualmodels] semantic retrieval. For each $y$ action label:
@@ 4.3 Kinematic Alignment via MIC and Contrastive Learning
Semantic similarity does not guarantee kinematic relevance. Moreover, $\mathcal{D}^{P}$ motions, designed for motion understanding through natural language, span longer time periods than atomic $\mathcal{D}^{T}$ actions designed for activity recognition, creating a structural disparity (see supplementary material for sequence-length analysis). We address both via the Mining In Context (MIC) module, which learns frame-level alignments for targeted window extraction.
@@ 4.3.1 MIC processing
    The tokens from both streams are fed into the MIC module, a bidirectional GRU with attention [bahdanau2016neuralmachinetranslationjointly, gru], producing context-aware latent vectors $z^{P}$ and $z^{\bar{T}}$.
@@ 4.3.2 Contrastive Alignment
MIC is trained to align $z^{\bar{T}_{i}}$ with $z^{P_{j}}$ from soft-positive pairs sharing the same target class $y$. Given a training batch of size $B$, we use the Soft Nearest Neighbors loss [pmlr-v2-salakhutdinov07a, frosst2019analyzingimprovingrepresentationssoft], which naturally accommodates the one-to-many soft positives:
$L_{con}=-\sum_{i\in B}\log\frac{\sum_{j\in B^{+}(i)}\exp(\text{sim}(z^{\bar{T}_{i}},z^{P_{j}})/\tau)}{\sum_{k\in B}\exp(\text{sim}(z^{\bar{T}_{i}},z^{P_{k}})/\tau)}$ (1)
where $B^{+}(i)$ contains batch indices $j$ whose $z^{P_{j}}$ was retrieved as a soft positive for the same class $y$ as $z^{\bar{T}_{i}}$. This enforces kinematic alignment within semantically-matched pairs via contrastive learning, while MIC’s attention enables subsequent window extraction.
@@ 4.4 Kinematic Mining
With MIC trained via contrastive loss, its attention mechanism learns to focus on kinematically relevant frames in $\mathbf{x}^{P}$ to minimize the enforced alignment loss with $\mathbf{x}^{\bar{T}}$. Let $\mathbf{a}^{P}=\{a^{P}_{k}\}_{k=1}^{n}$ denote attention weights over $\mathbf{x}^{P}$ of length $n$, and let $m$ be the length of $\mathbf{x}^{\bar{T}}$. We extract the most relevant contiguous subsequence (the "prior window") as the $m$-frames segment maximizing cumulative attention:
$\mathbf{x}^{P^{*}}=\underset{\{x^{P}(i),\ldots,x^{P}(i+m)\}\subseteq\mathbf{x}^{P}}{\arg\max}\sum_{k=i}^{i+m}a^{P}(k)$ (2)
The identified prior window $\mathbf{x}^{P^{*}}$ serves as a pseudo-labeled training sample for $G^{T}$ under the same target conditioning $c^{T}$ (Fig. 3).
@@ 4.5 Optimization of Multi-Objective Loss Function
We distill the teacher’s internal representations relative to the mined window into the student. Let $\mathbf{u}^{P}$ and $\mathbf{u}^{P^{*}}$ denote pre-projection feature sequences for $G^{P}$ and $G^{T}$, respectively. For a prior window starting at frame index $i$, the distillation loss is:
@@
$dww=\frac{1+\text{sim}(z^{\bar{T}},z^{P^{*}})}{2}$ (6)
where $z^{P^{*}}$ is computed with gradients detached, ensuring MIC remains trained exclusively by Eq. 1. The score dwwdww scales both reconstruction (Eq. 4) and distillation (Eq. 5) losses, giving higher weight to more reliable windows.
@@
The final optimization objective is a weighted combination of all components:

### v2

@@ Abstract
The acquisition cost for large, annotated motion datasets remains a critical bottleneck for skeletal-based Human Activity Recognition (HAR). Although Text-to-Motion (T2M) generative models offer a compelling, scalable source of synthetic data, their training objectives, which emphasize general artistic motion, and dataset structures fundamentally differ from HAR’s requirements for kinematically precise, class-discriminative actions. This disparity creates a significant domain gap, making generalist T2M models ill-equipped for generating motions suitable for HAR classifiers. To address this challenge, we propose KineMIC (Kinetic Mining In Context), a transfer learning framework for few-shot action synthesis. KineMIC adapts a T2M diffusion model to an HAR domain by hypothesizing that semantic correspondences in the text encoding space can provide soft supervision for kinematic distillation. We operationalize this via a kinetic mining strategy that leverages CLIP text embeddings to establish correspondences between sparse HAR labels and T2M source data. This process guides fine-tuning, transforming the generalist T2M backbone into a specialized few-shot Action-to-Motion generator. We validate KineMIC using HumanML3D as the source T2M dataset and a subset of NTU RGB+D 120 as the target HAR domain, randomly selecting just 10 samples per action class for training. Our approach generates significantly more coherent motions, providing a robust data augmentation source that delivers a +23.1% accuracy points improvement. Animated illustrations and supplementary materials are available at https://lucazzola.github.io/publications/kinemic.
@@ 1 Introduction
Human Activity Recognition (HAR) has become a cornerstone in a multitude of fields, including sports performance analysis, human-robot collaboration, and intelligent surveillance [6]. Skeletal-based HAR remains a fundamental modality, frequently employed due to its lightweight representation, robustness to environmental variations, and inherently privacy-preserving nature [38]. However, the performance of deep learning models for this task is fundamentally limited by the availability of large-scale, accurately annotated datasets. The acquisition and precise labeling of such high-quality, task-specific data is notoriously expensive and labor-intensive, creating a significant bottleneck that hampers progress, particularly in few-shot settings [6, 17]. This core challenge of data scarcity in HAR is what we aim to address.
@@
To mitigate this fundamental data bottleneck, much of modern few-shot recognition research focuses on applying strategies such as meta-learning and metric-learning frameworks directly to the action recognition model [37]. An alternative approach leverages generative models to create novel synthetic samples, thereby augmenting the small training set [12]. When it comes to motion synthesis, current research is dominated by Text-to-Motion (T2M) synthesis models, due to the growing interest in text as a conditioning modality. This interest has driven the creation and collection of large-scale T2M datasets [14, 24], which has, in turn, led to the development of powerful deep models [13, 14, 35] that can be employed as general motion priors for other generative tasks [30, 31, 34]. T2M priors are strategically appealing as a generative solution because the scalability of their annotation, using free-form, descriptive text, is significantly easier to achieve for general diversity than collecting high-volume, kinematically specific data for HAR.
@@
While the community moves towards developing foundation models for 3D humans [18, 36], the challenge of effectively exploiting such general T2M priors for specialized, downstream tasks like HAR remains underexplored. Our work concentrates on this challenge: adapting a general T2M prior to function as an Action-to-Motion (A2M) synthetic data generator for a specific target HAR domain. This transformation is non-trivial due to a significant domain gap characterized by two key factors. First, a semantic discrepancy exists where source T2M data uses descriptive text, while target HAR requires generation based on discrete action labels. Second, a kinematic gap exists between the broad, fluid motions of the source domain and the short, atomic motions required for HAR. The pre-trained T2M model, being a generalist, is therefore ill-equipped to meet requirements for reliable HAR classification.
@@
To address these challenges and bridge this domain gap for few-shot action synthesis, we propose KineMIC (Kinetic Mining In Context, Fig. 1). Our method employs teacher-student architecture, wherein a frozen teacher, pre-trained on a source T2M dataset, guides the fine-tuning of a student model on the limited target HAR domain. Our method starts by establishing a semantic correspondence between the sparse target action labels and the rich source textual descriptions through CLIP [28] text encoder. Secondly, relevant motion sub-sequences, extracted from the vast source dataset, are used to turn the general-purpose student into a specialized generator. The main contributions of this work are as follows:
@@
  • We introduce KineMIC, a teacher-student framework, which ultimately adapts a general T2M diffusion prior [35] to a specific HAR domain with minimal data, proving its effectiveness at improving HAR accuracy.
@@ 2.1 Skeletal-based HAR
Recognizing human activity from skeletal data is a pivotal research area in computer vision. Early approaches modeled motion as a time series using RNNs and LSTMs [8, 32]. A paradigm shift occurred with the introduction of Graph Convolutional Networks (GCNs) [38], which re-conceptualized the skeleton as a graph to model spatio-temporal dependencies. This led to rapid advancements using adaptive graph structures [33], refined GCNs [2, 3, 9, 22], and more recently, specialized architectures like 3D convolutional networks [10] and Transformers [7, 27]. Despite these advancements, the field still struggles with data scarcity, as the high capacity of modern models leads to severe overfitting in limited-data regimes. Our work addresses this bottleneck by proposing a novel synthesis framework leveraging deep T2M models, to create the necessary training data, supporting HAR applications in few-shot settings.
@@ 2.2 Generative Models for 3D Skeleton-based Motion
Realistic human motion synthesis evolved from early VAEs and GANs conditioned on discrete actions [15, 25] to highly expressive models. This progress was fueled by large motion capture datasets [14, 24] and richer conditioning signals, advancing from text [26] to even music [39] as a modality. The current state-of-the-art is dominated by Denoising Diffusion Probabilistic Models (DDPMs) [35, 40] and masked generative models [13], which achieve impressive generative fidelity. However, these T2M models, trained for general character animation, inherently lack the kinematic specificity required for downstream HAR tasks. Our proposed KineMIC framework addresses this domain gap by using a kinetic mining strategy to distill HAR-relevant knowledge from the T2M backbone, enabling its application as a specialized data generator.
@@ 2.3 Few-Shot HAR with Generative Models
The high cost of annotated motion data bottlenecks Few-Shot Human Activity Recognition (FSHAR). While classifier-side methods (e.g., metric/meta-learning) [37] are the most popular approach, they are inherently constrained by the limited kinematic diversity of the few-shot support set. A less-explored alternative is deep generative data augmentation [20, 23]. Most notably, recently Fukushi et al. [12] demonstrated this using a GAN with cross-domain regularization. To bypass GAN notorious training instability and explore a new avenue, we propose leveraging modern T2M diffusion models for few-shot motion synthesis. In direct contrast to [12], our approach employs a "semantics-first" matching strategy. By utilizing CLIP [28] text encoding space, we disentangle the synthesis process from the kinematic limitations of the few-shot set.
@@ 3 Problem Formulation
The core challenge we address is the adaptation of a pre-trained T2M generative model for data augmentation within the context of few-shot HAR. Our goal is to leverage the extensive kinematic knowledge contained in a large source (i.e. prior) domain to synthesize a high volume of diverse, class-specific motion sequences for a target domain, thereby enhancing the performance of a downstream HAR classifier. Let a skeletal motion sequence be defined as $\mathbf{x}=\{x(j)\in\mathbb{R}^{d}\}_{j=0}^{n-1}$, where $n$ is the number of frames and $d$ is the dimensionality of the pose representation. We consider two distinct domains:
@@ 4 Methodology
Figure 3: (Left) Core pipeline: dashed lines show detached gradients; dotted lines show weight sharing. MIC receives frame-wise tokens from teacher ($G^{P}$) and student ($G^{T}$) streams, producing latents $z^{P}$ and $z^{\bar{T}}$ for contrastive alignment and window mining. (Right) MIC module: attention-enhanced biGRU [1, 4] encoder aggregating frame-wise motion tokens into context-aware latents.
@@ 4.1 Teacher-Student Architecture
KineMIC builds on two identical MDM [35] models: a frozen teacher $G^{P}$ and a trainable student $G^{T}$. Both models retain MDM’s core structure and diffusion paradigm: predicting clean motion $\mathbf{x}_{0}$ (i.e. $\mathbf{x}$) from noised input $\mathbf{x}_{t}$ under a conditioning signal $c$. We refer to [35] for full diffusion process details. To preserve prior knowledge, $G^{T}$ is fine-tuned with Low-Rank Adaptation (LoRA) [16], which updates only a small fraction of weights through low-rank matrices while keeping the vast majority of pre-trained parameters frozen. Furthermore, while the student inherits the frozen text embedding layer used in T2M, we introduce a learnable action embedding (Fig. 2) that enables $G^{T}$ specialization to discrete HAR labels.
@@ 4.2 Semantic Alignment via Soft Positive Search
Target labels in $\mathcal{D}^{T}$ lack the descriptive richness of T2M captions in $\mathcal{D}^{P}$, creating a one-to-many mapping (single label $\rightarrow$ multiple text prompts). We bridge this via CLIP-based [28] semantic retrieval. For each $y$ action label:
@@ 4.3 Kinematic Alignment via MIC and Contrastive Learning
Semantic similarity does not guarantee kinematic relevance. Moreover, $\mathcal{D}^{P}$ motions, designed for motion understanding through natural language, span longer time periods than atomic $\mathcal{D}^{T}$ actions designed for activity recognition, creating a structural disparity (see supplementary material 2.4 for sequence-length analysis). We address both via the Mining In Context (MIC) module, which learns frame-level alignments for targeted window extraction.
@@ 4.3.1 MIC processing
    Tokens from both streams are fed into the MIC module, a bidirectional GRU with attention [1, 4], producing context-aware latent vectors $z^{P}$ and $z^{\bar{T}}$.
@@ 4.3.2 Contrastive Alignment
MIC is trained to align $z^{\bar{T}_{i}}$ with $z^{P_{j}}$ from soft-positive pairs sharing the same target class $y$. Given a training batch $B=\{1,\ldots,b\}$ of size $b$, we use the Soft Nearest Neighbors loss [29, 11], which naturally accommodates the one-to-many soft positives:
$L_{con}=-\frac{1}{b}\sum_{i\in B}\log\frac{\sum_{j\in B^{+}(i)}\exp(\text{sim}(z^{\bar{T}_{i}},z^{P_{j}})/\tau)}{\sum_{k\in B\setminus\{i\}}\exp(\text{sim}(z^{\bar{T}_{i}},z^{P_{k}})/\tau)}$ (1)
where $B^{+}(i)$ contains indices $j\in B\!\setminus\!\{i\}$ whose $z^{P_{j}}$ was retrieved as a soft positive for the same class $y$ as $z^{\bar{T}_{i}}$. This enforces kinematic alignment via a contrastive objective where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ temperature. This process allows MIC’s attention to enable subsequent window extraction.
@@ 4.4 Kinematic Mining
With MIC trained via contrastive loss, its attention mechanism learns to focus on kinematically relevant frames in $\mathbf{x}^{P}$ to minimize the enforced alignment loss with $\mathbf{x}^{\bar{T}}$. Let $\mathbf{a}^{P}=\{a^{P}(k)\}_{k=0}^{n-1}$ denote attention weights over $\mathbf{x}^{P}$ of length $n$, and let $m$ be the length of $\mathbf{x}^{\bar{T}}$. We extract the most relevant contiguous subsequence (the "prior window") as the $m$-frame segment maximizing cumulative attention:
$\mathbf{x}^{P^{*}}_{t}=\underset{\{x^{P}_{t}(i),\ldots,x^{P}_{t}(i+m-1)\}\subseteq\mathbf{x}^{P}_{t}}{\arg\max}\sum_{k=i}^{i+m-1}a^{P}(k)$ (2)
The identified prior window $\mathbf{x}^{P^{*}}_{t}$ serves as a pseudo-labeled training sample for $G^{T}$ under the same target conditioning $c^{T}$ (Fig. 3).
@@ 4.5 Optimization of Multi-Objective Loss Function
We distill the teacher’s internal representations relative to the mined window into the student. Let $\mathbf{u}^{P}$ and $\mathbf{u}^{P^{*}}$ denote pre-projection feature sequences of $\mathbf{x}^{P}_{t}$ and $\mathbf{x}^{P^{*}}_{t}$ processed by $G^{P}$ and $G^{T}$, respectively. For a prior window of length $m$ starting at frame index $i$, the distillation loss is:
@@
$dww=(1+\text{sim}(z^{\bar{T}},z^{P^{*}}))/2$ (6)
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity and $z^{P^{*}}$ is computed with gradients detached, ensuring MIC remains trained exclusively by Eq. 1. The score dwwdww scales both reconstruction (Eq. 4) and distillation (Eq. 5) losses, giving higher weight to more reliable prior windows.
@@
Final optimization follows a batch-averaged weighted objective: